OpenStudy (anonymous):

Insert grouping symbols so that the given expression has a value of 34. 14 + 10 – 6 • 8 – 3 A. (14 + 10) – 6 • (8 – 3) B. (14 + 10) – (6 • 8 – 3) C. (14 + 10) – (6 • 8) – 3 D. 14 + (10 – 6) • (8 – 3)

OpenStudy (anonymous):

d. 14+ (4) * (5) = 14 + 20 = 34

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem(: